# FullStack Bench: Evaluating LLMs as Full Stack Coders

### v3

@@ 4 Experiments
Qwen2.5-Coder-7B-Instruct 36.51 52.06 46.00 59.39 48.95 50.00 37.50 30.51 53.17 50.00 63.24 53.19 48.16
@@ 4.1 Experimental Setup
For open-sourced coders, we pull model checkpoints from Hugging Face [^7] https://huggingface.co/ and load the model with AutoTokenizer and AutoModelForCausalLM in the transformers package  [^8] https://pypi.org/project/transformers/. Further, we apply vLLM [Kwon et al., 2023] to accelerate the evaluation. We set top_k to 1, max_tokens to 2048, and keep all other settings as default. For those model versions not supported by vLLM, we call the default generate() method provided by the model and set top_k to 1, max_new_tokens to 1024. For the input prompt, we set the “You are a helpful AI assistant” as the system prompt and set each problem text in FullStack Bench as the user prompt. Then we apply the default chat template within the tokenizer and feed the tokens into the model.
We chat with the close-sourced models via API calls similarly, where top_k is set to 1 and max_new_tokens is set to 2048. The prompt template is kept the same. After model inference, the first code block with the corresponding programming language formatted in Markdown is extracted from the generated output. If no Markdown code block is detected, a heuristic approach is employed to identify and extract incomplete code snippets. The extracted code, combined with predefined test cases, is then used to synthesize the complete code, which is evaluated for correctness.
@@ 4.2 Results and Analysis
Qwen2.5-Coder-7B-Instruct 93.33 35.05 46.97 25.00 44.12 63.75 53.93 70.51 72.22 47.16 15.00 47.37 53.33 26.79 60.00 65.79 48.16

### v4

@@ 4 Experiments
Qwen2.5-Coder-7B-Instruct 38.60 53.23 39.00 63.20 49.65 44.56 37.50 33.46 46.83 55.00 63.24 54.26 47.95
@@ 4.1 Experimental Setup
For open-sourced coders, we pull model checkpoints from Hugging Face [^7] https://huggingface.co/ and load the model with vLLM [Kwon et al., 2023] to accelerate the evaluation. We set temperature to 0, max_completion_tokens to 2048, and keep all other settings as default. For the input prompt, we set each problem text in FullStack Bench as the user prompt, while leaving the system prompt as default for each model.
We chat with the close-sourced models via API calls similarly, where temperature is set to 0 and max_completion_tokens is set to 2048. The prompt template is kept the same. After model inference, the first code block with the corresponding programming language formatted in Markdown is extracted from the generated output. If no Markdown code block is detected, a heuristic approach is employed to identify and extract incomplete code snippets. The extracted code, combined with predefined test cases, is then used to synthesize the complete code, which is evaluated for correctness.
@@ 4.2 Results and Analysis
Qwen2.5-Coder-7B-Instruct 93.33 34.58 58.08 23.91 54.41 65.63 43.01 73.08 63.89 47.22 18.75 52.63 56.67 35.71 60.00 68.42 47.95